# TS-RAG: Retrieval-Augmented Generation based Time Series Foundation Models are Stronger Zero-Shot Forecaster

### v2

@@ 4.3.4 Effectiveness on Longer Forecasting Horizons
4.4 Case Studies
To further illustrate the effectiveness of TS-RAG, we conduct case studies on retrieval quality and its impact on forecasting performance compared to the baseline TSFM (Chronos-bolt). As shown in Figure 3, the top panel presents the input query and corresponding forecasting horizon from the Weather dataset, while the bottom panel displays the retrieved sequences. The retriever selects the most relevant historical sequences based on the embedding similarity of query time series. The retrieved sequences exhibit strong alignment with the input in terms of both trend and periodicity, indicating that the retrieval process effectively captures complex temporal patterns. Moreover, the forecasting horizons associated with the retrieved sequences closely align with the real future horizon, demonstrating that retrieved-augmented forecasting provides valuable external information to refine and explain the forecasts.
Figure 4 showcases a case study from the ETTm1 dataset. The top panel illustrates the input query, its corresponding forecasting horizon, and the future horizon of a retrieved sequence, while the bottom panel compares the forecasting results of the backbone TSFM and TS-RAG. The TSFM without RAG fails to capture a sudden trend shift effectively. By retrieving similar patterns from the retrieval knowledge base, TS-RAG successfully adapts to the trend, leading to more accurate forecasts. More case studies are shown in Appendix B and demonstrate how retrieval-augmented forecasting enhances robustness and interpretability across diverse real-world time series.
@@ 5 Conclusion
In this paper, we introduced TS-RAG, a novel retrieval-augmented forecasting framework designed to enhance the generalization and interpretability of Time Series Foundation Models (TSFMs) in zero-shot forecasting. By integrating retrieval-augmented generation (RAG) with a pretrained retrieval encoder and a Mixture-of-Experts (MoE) based augmentation module, TS-RAG effectively incorporates retrieved relevant patterns to improve forecasting accuracy in previously unseen domains. Extensive empirical evaluations on multiple benchmark datasets demonstrate that TS-RAG consistently outperforms existing TSFMs, validating its effectiveness in zero-shot forecasting while maintaining strong interpretability.
Looking ahead, we aim to 1) explore multimodal extensions of TS-RAG by integrating heterogeneous time series data, such as text data, to further enhance forecasting capabilities; 2) we plan to investigate optimization techniques for retrieval ranking in RAG, assessing whether more effective retrieval mechanisms can further boost zero-shot forecasting performance. We believe that TS-RAG establishes a strong foundation for retrieval-augmented time series forecasting, setting up a new frontier for robust and adaptable time series forecasting in dynamic and open world environments.
Impact Statement
@@ A.2 Baseline Introduction
  • Chronos-bolt (Ansari et al., 2024): Chronos-bolt is a subsequent version of Chronos, which can handle patch-based and uses decoder representations to generate quantile forecasts across multiple future steps, improving forecast accuracy over Chronos.
@@
  • MOMENT (Goswami et al., 2024): MOMENT uses a masking modeling technique for zero-shot forecasting by appending a lookback series with a mask that matches the length of the forecast. It involves pretraining a Transformer encoder model univariately on the ”Time Series Pile” datasets, which includes a wide variety of time series data.
@@
  • TTM (Ekambaram et al., 2024): TTM pre-trains a compact model based on the lightweight TSMixer architecture. It incorporates adaptive patching, diverse resolution sampling, and resolution prefix tuning to pretrain successfully on a small dataset.
@@
  • Moirai (Woo et al., 2024): Moirai pretrains the Transformer encoder on the “LOTSA” dataset, which includes27B time points, by masking the forecast horizon of each target channel and performing mask reconstruction.
@@
  • TimesFM (Das et al., 2023): TimesFM employs a decoder-style attention model and is pre-trained in a univariate manner on a large group of both real-world and synthetic datasets.
@@
  • Chronos (Ansari et al., 2024): Chronos is a probabilistic time series foundation model.Chronos tokenizes the input time series in a quantized manner and processes these tokens using the T5 model (Raffel et al., 2020). Chronos is trained on an extensive corpus of collected and synthetic time series data and has great generalization ability.
A.3 Details of Inference Datasets
@@
We experiment the zero-shot forecasting on the widely adopted Electricity Transformer Temperature (ETT) datasets (Zhou et al., 2021), Weather, Electricity (Wu et al., 2023), and Exchange Rate from  (Lai et al., 2018). ETT datasets are comprised of roughly two years of data from two locations in China. The data are further divided into four distinct datasets, each with different sampling rates: ETTh1 and ETTh2 are sampled hourly, and ETTm1 and ETTm2 are sampled every 15 minutes. Every ETT dataset includes six power load features and a target variable: the oil temperature. The Electricity dataset comprises records of electricity consumption from 321 customers and is measured with a 1-hour sampling rate. The Weather dataset contains one-year records from 21 meteorological stations located in Germany. The sampling rate for the Weather dataset is 10 minutes. The Exchange Rate dataset includes the daily exchange rates of eight foreign countries, including Australia, British, Canada, Switzerland, China, Japan, New Zealand, and Singapore ranging from 1990 to 2016.
A.4 Evaluation Metrics
@@
Appendix B Showcases
B.1 Case Studies on Retrieval Effectiveness
Figures 5 and  6 illustrate the retrieval performance of TS-RAG on the ETTh1 and ETTh2 datasets.The retrieval results demonstrate that TS-RAG effectively identifies historical patterns with strong structural similarity to the input, particularly in terms of periodicity and trend dynamics. In ETTh1, the retrieved sequences capture complex fluctuations and local variations, aligning well with the seasonal patterns of the input. Meanwhile, in ETTh2, where the time series exhibits smoother periodicity, the retrieved sequences show almost perfect alignment, indicating the presence of highly consistent cyclic behavior. These results suggest that retrieval augmentation enhances forecasting by leveraging historical patterns that closely match the current context, particularly in datasets with strong seasonal dependencies.
@@ B.1 Case Studies on Retrieval Effectiveness
B.2 Case Studies on Retrieval-Augmented Forecasting
Figures 7 and  8 showcase the impact of retrieval augmentation on forecasting accuracy in the Weather dataset. Figure 7 highlights a situation that the baseline TSFM struggles to capture a sudden trend shift, leading to a significant forecasting error. By incorporating retrieved forecasting horizons, TS-RAG successfully adapts to the trend change. Figure 8 demonstrates how retrieval augmentation enhances peak prediction. The standard TSFM underestimates the upcoming peak, whereas TS-RAG, guided by similar retrieved patterns, generates a more accurate forecast. These case studies illustrate how retrieval-augmented forecasting helps models better adapt to complex temporal patterns, improving robustness in real-world forecasting tasks.

### v3

@@ 5 Conclusion
In this paper, we introduced TS-RAG, a novel retrieval-augmented forecasting framework designed to enhance the generalization and interpretability of Time Series Foundation Models (TSFMs) in zero-shot forecasting. By integrating retrieval-augmented generation (RAG) with a pretrained retrieval encoder and an Adaptive Retrieval Mixer (ARM) augmentation module, TS-RAG effectively incorporates retrieved relevant patterns to improve forecasting accuracy in previously unseen domains. Extensive empirical evaluations on multiple benchmark datasets demonstrate that TS-RAG can consistently enhance the zero-shot forecasting performance of various TSFMs across diverse domains. Furthermore, we systematically explore the impact of different retrieval configurations, validating TS-RAG as a general and flexible framework for retrieval-augmented time series forecasting.
In summary, TS-RAG establishes a strong foundation for retrieval-augmented time series forecasting, setting up a new frontier for robust and adaptable time series forecasting in dynamic and open world environments. Looking ahead, we aim to 1) explore multimodal extensions of TS-RAG by integrating heterogeneous time series data, such as text data, to further enhance forecasting capabilities; 2) investigate optimization techniques for retrieval ranking in RAG, assessing whether more effective retrieval mechanisms can further boost zero-shot forecasting performance.
@@ A.2 TSFM Baseline Introduction
  • Chronos-bolt [25]: Chronos-bolt is a subsequent version of Chronos, which can handle patch-based and uses decoder representations to generate quantile forecasts across multiple future steps, improving forecast accuracy over Chronos.
@@
  • MOMENT [37]: MOMENT uses a masking modeling technique for zero-shot forecasting by appending a lookback series with a mask that matches the length of the forecast. It involves pretraining a Transformer encoder model univariately on the "Time Series Pile" datasets, which include a wide variety of time series data.
@@
  • TTM [24]: TTM pre-trains a compact model based on the lightweight TSMixer architecture. It incorporates adaptive patching, diverse resolution sampling, and resolution prefix tuning to pretrain successfully on a small dataset.
@@
  • Moirai [26]: Moirai pretrains the Transformer encoder on the “LOTSA” dataset, which includes 27B time points, by masking the forecast horizon of each target channel and performing mask reconstruction.
@@
  • TimesFM [23]: TimesFM employs a decoder-style attention model and is pre-trained in a univariate manner on a large group of both real-world and synthetic datasets.
@@
  • Chronos [25]: Chronos is a probabilistic time series foundation model.Chronos tokenizes the input time series in a quantized manner and processes these tokens using the T5 model [49]. Chronos is trained on an extensive corpus of collected and synthetic time series data and has great generalization ability.
  • Time-MOE [41]: Time-MoE consists of a family of decoder-only time series foundation models with a mixture-of-experts architecture, designed to operate in an auto-regressive manner, enabling universal forecasting with arbitrary prediction horizons and long context lengths.
A.3 RAF Implementation
Following the original RAF paper, we implement RAF using the Chronos-Base model as the backbone. To ensure a fair comparison, we align both the test set and the retrieval knowledge base with those used in TS-RAG. While the original RAF paper adopts WQL and MASE as evaluation metrics, we report MAE and MSE to maintain consistency with TS-RAG’s evaluation protocol.
For the retrieval process, we retrieve the top-10 most relevant sequences, consistent with the TS-RAG setting. Additionally, we implement an RAF variant using the Chronos-Bolt model as the backbone to fully align model configurations. These adjustments ensure that any observed performance differences come solely from the retrieval and augmentation mechanisms.
A.4 Details of Inference Datasets
@@
We experiment the zero-shot forecasting on the widely adopted Electricity Transformer Temperature (ETT) datasets [10], Weather, Electricity [50], and Exchange Rate from  [51]. ETT datasets are comprised of roughly two years of data from two locations in China. The data are further divided into four distinct datasets, each with different sampling rates: ETTh1 and ETTh2 are sampled hourly, and ETTm1 and ETTm2 are sampled every 15 minutes. Every ETT dataset includes six power load features and a target variable: the oil temperature. The Electricity dataset comprises records of electricity consumption from 321 customers and is measured with a 1-hour sampling rate. The Weather dataset contains one-year records from 21 meteorological stations located in Germany. The sampling rate for the Weather dataset is 10 minutes. The Exchange Rate dataset includes the daily exchange rates of eight foreign countries, including Australia, Britain, Canada, Switzerland, China, Japan, New Zealand, and Singapore, ranging from 1990 to 2016.
A.5 Evaluation Metrics
@@
A.6 Efficiency Analysis
Training Efficiency
TS-RAG is designed for efficient adaptation. It freezes the TSFM backbone and trains only a lightweight augmentation module (ARM + projection), reducing trainable parameters and improving stability. Retrieval indices are precomputed to avoid redundancy. Training on 26M pairs with 20,000 steps takes around 1 hour on a single NVIDIA A6000 GPU, using TF32 precision and applying dropout.
Inference Efficiency
At inference time, TS-RAG introduces a top-$k$ retrieval step. We use FAISS to perform fast nearest-neighbor search over the retrieval knowledge base. On ETTh, retrieval adds 9.2 ms latency per query; the ARM-augmented forward pass adds 0.44 ms. Total inference time is 9.62 ms/query. This overhead is minor and offset by consistent gains (e.g., --6.84% MSE on Exchange rate), making TS-RAG suitable for real-time use. Notably, compared to RAF, which requires 3.47 seconds per query, TS-RAG achieves over 360$\times$×faster inference speed, highlighting its superior efficiency and practicality.
Appendix B Additional Results
B.1 Generalization across Backbones
To verify the generalization ability of TS-RAG, we further evaluate it on the MOMENT [37] backbone, in addition to the Chronos-bolt results reported in the main text. The same retrieval-augmented framework is applied without modifying the backbone architecture or training recipes, and all retrieval settings follow those used with Chronos-bolt for fair comparison. Note that to enable the MOMENT model to perform zero-shot long-term forecasting, we pretrain a prediction head on the same pretraining data as TS-RAG. Table 4 reports the results on both backbones. TS-RAG consistently improves zero-shot forecasting performance regardless of the backbone, demonstrating its plug-and-play nature.
Table 4: Zero-shot forecasting results of TS-RAG across different backbones (Chronos-bolt and Moment). Both MSE and MAE are reported. The best results are highlighted in bold.
Backbone Methods ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange Average
Chronos-bolt MSE 0.3616 0.2517 0.3109 0.1487 0.1525 0.1132 0.0673 0.2008
MAE 0.3650 0.2992 0.3185 0.2236 0.1825 0.2004 0.1780 0.2525
TS-RAGChronos-bolt MSE 0.3557 0.2451 0.2906 0.1466 0.1454 0.1120 0.0627 0.1940
MAE 0.3624 0.2982 0.3114 0.2231 0.1771 0.2002 0.1718 0.2492
Moment MSE 0.3920 0.2742 0.3506 0.1703 0.1801 0.1967 0.0979 0.2374
MAE 0.4110 0.3327 0.3824 0.2579 0.2384 0.3028 0.2059 0.3044
TS-RAGMoment MSE 0.3823 0.2511 0.3325 0.1552 0.1604 0.1920 0.0775 0.2216
MAE 0.4072 0.3220 0.3738 0.2474 0.2212 0.2994 0.1972 0.2955
B.2 Ablation Study (Extended Results)
In this section, we present the extended results of the ablation studies reported in the main text. Due to space constraints, some detailed results were omitted or summarized in the main paper. Here, we provide the complete results to facilitate a more comprehensive understanding and reproducibility.
Table 5: Full Long-term zero-shot forecasting results across various TSFMs and TS-RAG. Both MSE and MAE are reported. The best results are highlighted in bold and second-best results are underlined.
Backbone Methods ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange Average
TS-RAGChronos-bolt MSE 0.3557 0.2451 0.2906 0.1466 0.1454 0.1120 0.0627 0.1940
MAE 0.3624 0.2982 0.3114 0.2231 0.1771 0.2002 0.1718 0.2492
Chronos-boltB MSE 0.3616 0.2517 0.3109 0.1487 0.1525 0.1132 0.0673 0.2008
MAE 0.3650 0.2992 0.3185 0.2236 0.1825 0.2004 0.1780 0.2525
MOMENT MSE 0.3920 0.2742 0.3506 0.1703 0.1801 0.1967 0.0979 0.2374
MAE 0.4110 0.3327 0.3824 0.2579 0.2384 0.3028 0.2059 0.3044
TTMB MSE 0.3619 0.2531 0.3152 0.1511 0.1543 0.1715 0.0657 0.2104
MAE 0.3710 0.3032 0.3248 0.2405 0.1893 0.2643 0.1725 0.2665
MoiraiB MSE 0.3686 0.2547 0.5399 0.1958 0.1711 0.1832 0.0663 0.2542
MAE 0.3835 0.3053 0.4322 0.2687 0.1912 0.2814 0.1720 0.2906
TimesFM MSE 0.4254 0.2894 0.3321 0.1703 0.0695
MAE 0.3825 0.3233 0.3326 0.2552 0.1802
ChronosB MSE 0.4217 0.2659 0.3935 0.1663 0.1897 0.1460 0.0831 0.2380
MAE 0.3806 0.3136 0.3695 0.2522 0.2107 0.2237 0.1879 0.2769
Time-MoE MSE 0.3623 0.2521 0.3213 0.1565 0.1490 0.1137 0.0851 0.2057
MAE 0.3669 0.3224 0.3340 0.2540 0.1844 0.2026 0.2056 0.2671
Table 6: Long-term zero-shot forecasting results with different retrieval knowledge bases. The best results are highlighted in bold, and the second-best results are underlined.
Knowledge Base ETTh1 ETTh2 ETTm1 ETTm2
MSE MAE MSE MAE MSE MAE MSE MAE
Cross-Domain 0.3601 0.3667 0.2466 0.2999 0.3046 0.3240 0.1496 0.2279
Distribution-Shift 0.3586 0.3647 0.2453 0.2996 0.2993 0.3179 0.1502 0.2271
Multi-Domain 0.3564 0.3633 0.2432 0.2973 0.2971 0.3157 0.1513 0.2277
In-Domain 0.3557 0.3624 0.2451 0.2982 0.2906 0.3114 0.1466 0.2231
Table 7: TS-RAG experiment results using different augmentation techniques.
Method Metric ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange Average
TS-RAGARM MSE 0.3557 0.2451 0.2906 0.1466 0.1454 0.1120 0.0627 0.1940
MAE 0.3624 0.2982 0.3114 0.2231 0.1771 0.2002 0.1718 0.2492
TS-RAGGate MSE 0.3575 0.2498 0.3041 0.1473 0.1501 0.1126 0.0663 0.1982
MAE 0.3640 0.2988 0.3154 0.2235 0.1815 0.2005 0.1768 0.2515
Chronos-bolt MSE 0.3616 0.2517 0.3109 0.1487 0.1525 0.1132 0.0673 0.2008
MAE 0.3650 0.2992 0.3185 0.2236 0.1825 0.2004 0.1780 0.2525
Table 8: Zero-shot forecasting results for extended forecasting horizons (MSE).
Horizon Methods ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange Average
96 w/o RAG 0.3859 0.2899 0.3323 0.1779 0.1777 0.1242 0.0993 0.2267
TS-RAG 0.3772 0.2812 0.3141 0.1753 0.1697 0.1226 0.0927 0.2189
192 w/o RAG 0.4446 0.3603 0.3838 0.2515 0.2244 0.1428 0.1926 0.2857
TS-RAG 0.4306 0.3474 0.3688 0.2462 0.2172 0.1413 0.1831 0.2764
336 w/o RAG 0.4850 0.4045 0.4374 0.3177 0.2838 0.1613 0.3437 0.3476
TS-RAG 0.4650 0.3839 0.4153 0.3115 0.2819 0.1593 0.3157 0.3347
720 w/o RAG 0.4841 0.4143 0.5285 0.4162 0.3673 0.2069 0.8100 0.4610
TS-RAG 0.4703 0.4017 0.4935 0.4164 0.3703 0.2050 0.6968 0.4363
B.3 Sensitivity to the Number of Retrieved Sequences
Figure 5: Parameter sensitivity to the number of retrieved sequences (k) on seven zero-shot evaluation datasets.
The impact of varying the number of retrieved sequences ($k$) on forecasting performance is illustrated in Figure 5. The x-axis represents the number of retrieved sequences, while the y-axis shows the corresponding Mean Squared Error (MSE). Across all datasets, increasing $k$ initially leads to a significant decrease in MSE, demonstrating that incorporating additional retrieved sequences helps refine predictions by leveraging retrieved patterns. However, beyond a certain threshold, the improvement plateaus and even decreases slightly in some datasets, indicating diminishing returns as $k$ increases.
Dataset-specific trends further reveal differences in sensitivity to $k$. For instance, ETTm1 and ETTm2 exhibit the most pronounced improvement as $k$ increases, with MSE rapidly declining before stabilizing. This suggests that these datasets benefit significantly from retrieval-augmented inference, likely due to strong temporal dependencies in their historical patterns. ETTh1 and ETTh2 show a similar trend but with a smaller overall reduction in MSE, indicating that while retrieval is beneficial, these datasets may already contain strong intrinsic signals, making additional augmentation less impactful. The Weather, Electricity, and Exchange Rate datasets display a steady decline in MSE with $k$ increasing, but the improvement becomes marginal as $k$ increases further, suggesting that a moderate number of retrieved sequences is sufficient.
Lookback Length Metric ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange Average
64 MSE 0.3540 0.2432 0.3114 0.1502 0.1491 0.1132 0.0678 0.1984
128 MSE 0.3572 0.2415 0.2935 0.1494 0.1526 0.1125 0.0674 0.1963
256 MSE 0.3539 0.2409 0.3195 0.1518 0.1518 0.1130 0.0662 0.1996
512 MSE 0.3557 0.2451 0.2906 0.1466 0.1454 0.1120 0.0627 0.1940
B.4 Effect of Retrieval Lookback Length
Table 9 presents the effect of different retrieval lookback lengths on zero-shot forecasting performance. Given an input sequence of length 512, we explore different retrieval configurations by using only the last 64, 128 or 256 time steps, or the full 512 time steps for retrieval. Across all datasets, longer retrieval lookback windows (256 or 512) yield relatively better performance, suggesting that incorporating a more extended historical context helps retrieve more relevant sequences. This finding demonstrates that retrieving from longer historical sequences generally improves the quality of retrieved sequences, leading to greater forecasting accuracy. However, in some cases, longer is not always better, indicating that excessive retrieval windows may introduce noise or irrelevant information. This suggests the potential for adaptive retrieval mechanisms that allow the retriever to dynamically determine the most suitable retrieval lookback length for each instance.
Table 10: Performance comparison of TS-RAGChronos-bolt using different Retriever Encoders (Chronos, TTM, and MOMENT). Both MSE and MAE are reported.
Retriever Encoder ETTh1 ETTh2 ETTm1 ETTm2
MSE MAE MSE MAE MSE MAE MSE MAE
w/o RAG 0.3616 0.3650 0.2517 0.2992 0.3109 0.3185 0.1487 0.2236
Chronos 0.3557 0.3624 0.2451 0.2982 0.2906 0.3114 0.1466 0.2231
TTM 0.3556 0.3625 0.2465 0.2993 0.2906 0.3121 0.1452 0.2217
MOMENT 0.3553 0.3635 0.2437 0.2981 0.2967 0.3164 0.1465 0.2234
B.5 Impact of Retriever Encoder Choice
To further investigate the impact of retriever encoder choice, we conduct an ablation study using pretrained encoders from two additional TSFMs, namely TTM and MOMENT, as alternatives to the Chronos encoder used in our main experiments. TTM employs an MLP-Mixer-like architecture, while MOMENT is based on a Transformer encoder. The Chronos encoder is built on a T5 architecture. All experiments follow the same retrieval-augmented setup for fair comparison.
Table 10 reports the results. We observe that the performance across the three encoders is largely comparable on all datasets, and no encoder consistently outperforms the others. This suggests that the choice of retriever encoder architecture (among existing TSFMs) has limited impact on the overall performance of TS-RAG.
B.6 Comparion between TS-RAG and RAF
Table 11: Full results for zero-shot forecasting comparison between TS-RAG and RAF.
Method Metric ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange Average
RAFChronos MSE 0.4212 0.2661 0.3927 0.1659 0.1726 0.1377 0.0666 0.2318
MAE 0.3800 0.3137 0.3695 0.2521 0.2048 0.2189 0.1774 0.2738
RAFChronos-bolt MSE 0.3660 0.2524 0.3058 0.1477 0.1780 0.1185 0.0632 0.2045
MAE 0.3710 0.3046 0.3285 0.2281 0.2065 0.2114 0.1725 0.2604
TS-RAGChronos-bolt MSE 0.3557 0.2451 0.2906 0.1466 0.1454 0.1120 0.0627 0.1940
MAE 0.3624 0.2982 0.3114 0.2231 0.1771 0.2002 0.1718 0.2492
As shown in Table 11, the original RAF implementation (RAFChronos) exhibits significantly inferior performance compared to TS-RAG across all benchmarks. Even when upgrading the backbone of RAF to Chronos-bolt (RAFChronos-bolt), TS-RAGChronos-bolt still consistently outperforms RAF on both MSE and MAE metrics. These results demonstrate the effectiveness of TS-RAG’s overall design, which integrates both an optimized retrieval process and an adaptive augmentation module to enhance zero-shot forecasting performance.
Appendix C Showcases
@@
C.1 Case Studies on Retrieval Effectiveness
Figures 6 and  7 illustrate the retrieval performance of TS-RAG on the ETTh1 and ETTh2 datasets. The retrieval results demonstrate that TS-RAG effectively identifies historical patterns with strong structural similarity to the input, particularly in terms of periodicity and trend dynamics. In ETTh1, the retrieved sequences capture complex fluctuations and local variations, aligning well with the seasonal patterns of the input. Meanwhile, in ETTh2, where the time series exhibits smoother periodicity, the retrieved sequences show almost perfect alignment, indicating the presence of highly consistent cyclic behavior. These results suggest that retrieval augmentation enhances forecasting by leveraging time series patterns that closely match the current context, particularly in datasets with strong seasonal dependencies.
C.2 Case Studies on Retrieval-Augmented Forecasting
Figures 8 and  9 showcase the impact of retrieval augmentation on forecasting accuracy in the Weather dataset. Figure 8 highlights a situation where the baseline TSFM struggles to capture a sudden trend shift, leading to a significant forecasting error. By incorporating retrieved forecasting horizons, TS-RAG successfully adapts to the trend change. Figure 9 demonstrates how retrieval augmentation enhances peak prediction. The standard TSFM underestimates the upcoming peak, whereas TS-RAG, guided by similar retrieved patterns, generates a more accurate forecast. These case studies illustrate how retrieval-augmented forecasting helps models better adapt to complex temporal patterns, improving robustness in real-world forecasting tasks.